# Demystifying Cost-Efficiency in LLM Serving over Heterogeneous GPUs

### v1

@@ 1 Introduction
Currently, predominant practices utilize homogeneous GPU resources to deploy LLMs and serve the incoming requests (Li et al., 2023; Kwon et al., 2023; Agrawal et al., 2024). However, with the broadening application domains, serving LLMs is facing progressively varying request patterns, driving the serving workloads dynamic and diverse— a phenomenon referred to as workload heterogeneity (Sun et al., 2024; Zhao et al., 2024c). This contradiction makes the use of homogeneous GPU resources unsuitable.
@@ 2 Background
Workload heterogeneity. LLMs are designed to support a diverse range of applications, and the inference workloads associated with these applications often exhibit heterogeneity in input and output token lengths, which is called workload heterogeneity. Different workload types exhibit varying characteristics in terms of compute and memory demands. For example, requests from the WildGPT dataset (Zhao et al., 2024b), with average input and output token lengths of 496 and 510 respectively (classified as short input and long output), typically require more memory resources to handle the memory-bound decoding phase. Conversely, requests from the Azure-Trace dataset (Patel et al., 2024; Azure, 2024), with average input and output token lengths of 2455 and 18 respectively (classified as long input and short output), generally demand more compute resources to manage the compute-bound prefill phase. Therefore, appropriately allocating resources based on workload demands is critical for optimal performance.
Heterogeneous LLM serving. Recent research has explored various approaches for deploying LLM serving in heterogeneous GPU environments to achieve cost-efficient solutions (Jiang et al., 2023; Miao et al., 2024; Griggs et al., 2024; Zhao et al., 2024a; Mei et al., 2024; Borzunov et al., 2022). HexGen introduces asymmetric partitioning and advanced scheduling techniques to deploy generative inference in decentralized and heterogeneous settings. Me’lange frames GPU allocation as a cost-aware bin-packing problem, optimizing cost efficiency for LLM services by effectively leveraging heterogeneous GPU types. Helix formulates the problem of heterogeneous GPU and network connection optimization as a max-flow problem, utilizing mixed-integer linear programming to determine the optimal model deployment. However, existing works typically optimize performance within a predefined heterogeneous cluster, and fail to consider GPU availability and user-defined budget constraints on cloud platforms. In addition, they are generally unaware of the workload heterogeneity, and only consider uniform workload assignment.
@@ 3 Observation and Opportunity
[…]
@@
Benchmark settings. We subsample nine workload types from the ShareGPT (Zheng et al., 2023), WildGPT (Zhao et al., 2024b), and Azure-Trace datasets (Patel et al., 2024). These workloads are characterized by average input token lengths of $\{2455,824,496\}$ and output token lengths of $\{510,253,18\}$. Each combination reflects distinct workload characteristics. For example, $\{2455,18\}$ (long input, short output) represents compute-intensive workloads, while $\{496,510\}$ (short input, long output) represents memory-intensive workloads. Based on these workload types, we evaluate two models, Llama3-8B and Llama3-70B, on six commonly used cloud GPUs (A6000, A40, L40, A100, H100, and 4090) with different deployment configurations. The benchmarking metrics include request throughput per unit cost (i.e., throughput divided by GPU cost) and the total cost associated with various latency percentiles (e.g., p5, p10, p15, $\dots$, p95, p100). The total cost for each latency percentile is calculated by multiplying the latency time by the GPU cost. These metrics serve as indicators of cost efficiency. The GPU costs are demonstrated in Table 1.
@@ 4.3 MILP Formulation
Configurations. We consider a set $\mathcal{C}$ of feasible configurations (i.e., deployment configurations). Each configuration $c\in\mathcal{C}$ represents the serving plan for a single model replica, which is characterized by $(v_{c},s_{c},o_{c},h_{c,w})$: (i) A vector $v_{c}=\{d_{n}(c)\}_{n=1}^{N}$ indicating exactly how many GPUs of each type $n$ are used in configuration $c$. (ii) An array $s_{c}=\{t_{1},t_{2},\dots,t_{S}\}$ indicating the parallelism strategy used in configuration $c$. The array length $S$ represents the total number of pipeline stages, and the element $t_{s}$ represents the TP degree of the $s$-th stage. The summation of all $t_{s}$ should be equal to the total GPU count of configuration $c$, i.e., $\sum_{s=1}^{S}t_{s}=\sum_{n=1}^{N}d_{n}(c)$. (iii) A cost $o_{c}=\sum_{n=1}^{N}(d_{n}(c)\times p_{n})$ indicating the total price required for configuration $c$. (iv) A throughput $h_{c,w}$ indicating the rate at which configuration $c$ process workload type $w$, which is obtained through a one-time profiling. By optimizing the configurations, we can obtain the GPU composition and deployment configurations mentioned in §4.1.
Workloads and assignment. Let there be $W$ workload types, indexed by $w\in\{1,2,\dots,W\}$. Each workload $w$ must be fully served (i.e., 100% coverage). We allow fractional assignment: a fraction $x_{c,w}\in[0,1]$ of workload $w$ may be processed by configuration $c$. Concretely, $\sum_{c\in\mathcal{C}}x_{c,m}=1,\forall\,m=1,2,\dots,M$. We also introduce an integer variable $y_{c}\in\{0,1,2,\dots\}$ indicating how many copies of configuration $c$ are chosen (activated). If $y_{c}=0$, then $x_{c,w}$ must be zero for all $w$. By co-optimizing the workload assignment (the fractions $x_{c,w}$) with the activated configurations ($y_{c}$), we determine the final workload assignment as described in §4.1.
@@
Optimization objective. We define a makespan variable $T\geq 0$ to represent the overall completion time. For a configuration $c$, if it is instantiated $y_{c}$ times and processes fractions $x_{c,w}$ of workload $w$, each replica provides a throughput of $h_{c,w}$. Consequently, the total effective throughput for $c$ is $y_{c}\times h_{c,w}$, and the time required on $c$ is given by $T_{c}=\sum_{w=1}^{W}\frac{x_{c,w}}{y_{c}\cdot h_{c,w}}.$ Since all chosen configurations run in parallel, the system completes once the slowest configuration finishes. Thus, we have $T_{c}\leq T$ for all $c\in\mathcal{C}$. Our optimization objective is to minimize $T$.
@@ 5.2 Experimental Results
We also compare our method with the state-of-the-art heterogeneous serving framework, HexGen (Jiang et al., 2023). Since HexGen schedules workloads based on a fixed GPU composition, we evaluate it using two setups: (i) a uniform composition, where six GPU types are evenly allocated within the budget, and (ii) the optimal composition used by our method. As shown in Figure 7, HexGen with a uniform composition suffers up to 35% and on average 29% performance degradation due to suboptimal GPU allocation. Even with the optimal composition, our method outperforms HexGen by up to 18% and on average 14%, benefiting from workload-aware scheduling.
@@ Appendix E Extend to Multiple LLM serving
s.t. $\forall m:\begin{cases}\sum_{c\in\mathcal{C}_{m}}x_{c,w,m}=1,\,\forall\,w\in W_{m},\\[4.0pt] \sum_{w\in W_{m}}\frac{x_{c,w,m}}{y_{c,m}\cdot h_{c,w,m}}\leq T,\,\forall\,c\in\mathcal{C}_{m},\\[4.0pt] x_{c,w,m}\leq y_{c,m},\,\forall\,c\in\mathcal{C}_{m},\,\forall\,w\in W_{m},\end{cases}$ (9)

### v2

@@ 1 Introduction
Currently, predominant practices utilize homogeneous GPU resources to deploy LLMs and serve the incoming requests (Li et al., 2023; Kwon et al., 2023; Agrawal et al., 2024b). However, with the broadening application domains, serving LLMs is facing progressively varying request patterns, driving the serving workloads dynamic and diverse— a phenomenon referred to as workload heterogeneity (Sun et al., 2024; Zhao et al., 2024b). This contradiction makes the use of homogeneous GPU resources unsuitable.
@@ 2 Background
Workload heterogeneity. LLMs are designed to support a diverse range of applications, and the inference workloads associated with these applications often exhibit heterogeneity in input and output token lengths, which is called workload heterogeneity. Different workload types exhibit varying characteristics in terms of compute and memory demands. For example, requests from the WildGPT dataset (Zhao et al., ), with average input and output token lengths of 496 and 510 respectively (classified as short input and long output), typically require more memory resources to handle the memory-bound decoding phase. Conversely, requests from the Azure-Trace dataset (Patel et al., 2024; Azure, 2024), with average input and output token lengths of 2455 and 18 respectively (classified as long input and short output), generally demand more compute resources to manage the compute-bound prefill phase. Therefore, appropriately allocating resources based on workload demands is critical for optimal performance.
Heterogeneous LLM serving. Recent research has explored various approaches for deploying LLM serving in heterogeneous GPU environments to achieve cost-efficient solutions (JIANG et al., ; Miao et al., 2024; Griggs et al., 2024; Zhao et al., 2024a; Mei et al., 2024; Borzunov et al., 2023; Yan et al., 2024). HexGen introduces asymmetric partitioning and advanced scheduling techniques to deploy generative inference in decentralized and heterogeneous settings. Me’lange frames GPU allocation as a cost-aware bin-packing problem, optimizing cost efficiency for LLM services by effectively leveraging heterogeneous GPU types. Helix formulates the problem of heterogeneous GPU and network connection optimization as a max-flow problem, utilizing mixed-integer linear programming to determine the optimal model deployment. However, existing works typically optimize performance within a predefined heterogeneous cluster, and fail to consider GPU availability and user-defined budget constraints on cloud platforms. In addition, they are generally unaware of the workload heterogeneity, and only consider uniform workload assignment.
LLM serving optimization. There are several related works focusing on the optimization of LLM serving (Li et al., 2023; Yu et al., 2022; Kwon et al., 2023). QLM (Patke et al., 2024) focuses on SLO-aware serving and multi-node optimizations by refining request ordering; SarathiServe (Agrawal et al., 2024b) optimizes batching through prefill chunking to mitigate interference between the prefill and decoding stages; and Vidur (Agrawal et al., 2024a) develops an accurate simulator for deployment tuning. In contrast, our method is dedicated to achieving heterogeneous, cost-efficient serving in cloud environments.
@@ 3 Observation and Opportunity
Figure 3: Benchmarked results for Llama3-70B model with different GPU types on different workload types. The left three columns represent the throughput results, x-axis represents different GPU types, y-axis represents throughput per unit price (i.e., throughput divided by GPU cost). The right three columns represent the latency results, x-axis represents the P5-100 latency results (P5-P100 Latencies means from left to right, the x sticks represent P5 Latency, P10 Latency, P15 Latency …), y-axis represents total price (i.e., each latency time multiplied by GPU cost). Results for Llama3-8B model are demonstrated in Appendix A.
@@
Benchmark settings. We subsample nine workload types from the ShareGPT (Zheng et al., ), WildGPT (Zhao et al., ), and Azure-Trace datasets (Patel et al., 2024). These workloads are characterized by average input token lengths of $\{2455,824,496\}$ and output token lengths of $\{510,253,18\}$. Each combination reflects distinct workload characteristics. For example, $\{2455,18\}$ (long input, short output) represents compute-intensive workloads, while $\{496,510\}$ (short input, long output) represents memory-intensive workloads. Based on these workload types, we evaluate two models, Llama3-8B and Llama3-70B, on six commonly used cloud GPUs (A6000, A40, L40, A100, H100, and 4090) with different deployment configurations. The benchmarking metrics include request throughput per unit cost (i.e., throughput divided by GPU cost) and the total cost associated with various latency percentiles (e.g., p5, p10, p15, $\dots$, p95, p100). The total cost for each latency percentile is calculated by multiplying the latency time by the GPU cost. These metrics serve as indicators of cost efficiency. The GPU costs are demonstrated in Table 1.
@@ 4.3 MILP Formulation
$f_{w}$ total requests of workload $w$
Configurations. We consider a set $\mathcal{C}$ of feasible configurations (i.e., deployment configurations). Each configuration $c\in\mathcal{C}$ represents the serving plan for a single model replica, which is characterized by $(v_{c},s_{c},o_{c},h_{c,w})$: (i) A vector $v_{c}=\{d_{n}(c)\}_{n=1}^{N}$ indicating exactly how many GPUs of each type $n$ are used in configuration $c$. (ii) An array $s_{c}=\{t_{1},t_{2},\dots,t_{S}\}$ indicating the parallelism strategy used in configuration $c$. The array length $S$ represents the total number of pipeline stages, and the element $t_{s}$ represents the TP degree of the $s$-th stage. The summation of all $t_{s}$ should be equal to the total GPU count of configuration $c$, i.e., $\sum_{s=1}^{S}t_{s}=\sum_{n=1}^{N}d_{n}(c)$. (iii) A cost $o_{c}=\sum_{n=1}^{N}(d_{n}(c)\times p_{n})$ indicating the total price required for configuration $c$. (iv) A throughput $h_{c,w}$ indicating the rate at which configuration $c$ process workload type $w$, which is obtained through a one-time profiling  [^1] The detailed procedure of the one-time profiling is demonstrated in Appendix L.. By optimizing the configurations, we can obtain the GPU composition and deployment configurations mentioned in §4.1.
Workloads and assignment. Let there be $W$ workload types, indexed by $w\in\{1,2,\dots,W\}$. Each workload $w$ must be fully served (i.e., 100% coverage). We allow fractional assignment: a fraction $x_{c,w}\in[0,1]$ of workload $w$ may be processed by configuration $c$. Concretely, $\sum_{c\in\mathcal{C}}x_{c,w}=1,\forall\,w=1,2,\dots,W$. We also introduce an integer variable $y_{c}\in\{0,1,2,\dots\}$ indicating how many copies of configuration $c$ are chosen (activated). If $y_{c}=0$, then $x_{c,w}$ must be zero for all $w$. By co-optimizing the workload assignment (the fractions $x_{c,w}$) with the activated configurations ($y_{c}$), we determine the final workload assignment as described in §4.1.
@@
Optimization objective. We define a makespan variable $T\geq 0$ to represent the overall completion time. For a configuration $c$, if it is instantiated $y_{c}$ times and processes fractions $x_{c,w}$ of workload $w$, each replica provides a throughput of $h_{c,w}$. Let $f_{w}$ be the total number of requests for workload $w$. Consequently, the total effective throughput for $c$ is $y_{c}\times h_{c,w}$, and the time required on $c$ is given by $T_{c}=\sum_{w=1}^{W}\frac{x_{c,w}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\cdot f_{w}}}{y_{c}\cdot h_{c,w}}.$ Since all chosen configurations run in parallel, the system completes once the slowest configuration finishes. Thus, we have $T_{c}\leq T$ for all $c\in\mathcal{C}$. Our optimization objective is to minimize $T$.
@@ 5.2 Experimental Results
Comparison with HexGen. We also compare our method with the state-of-the-art heterogeneous serving framework, HexGen (JIANG et al., ). Since HexGen schedules workloads based on a fixed GPU composition, we evaluate it using two setups: (i) a uniform composition, where six GPU types are evenly allocated within the budget, and (ii) the optimal composition used by our method. As shown in Figure 7, HexGen with a uniform composition suffers up to 35% and on average 29% performance degradation due to suboptimal GPU allocation. Even with the optimal composition, our method outperforms HexGen by up to 18% and on average 14%, benefiting from workload-aware scheduling.
Compare with Helix. We conduct additional experiments comparing our system with Helix (Mei et al., 2024). Specifically, we compare our method against Helix’s single-cluster case (the optimal case reported in their paper) under a price budget of \$15 per hour on the AzureTrace dataset (i.e., trace 2). While Helix optimizes heterogeneous LLM serving using max-flow and MILP scheduling, our method explicitly considers workload heterogeneity and GPU composition optimization, resulting in greater cost efficiency. Experimental results show that our system outperforms Helix by 25–35%.
Table 3: Performance comparison of Helix and our method.
Llama-30B Llama3-70B
Helix 8.49 req/s 5.72 req/s
Ours 11.43 req/s (35$\uparrow$) 7.13 req/s (25$\uparrow$)
@@
Additional discussion and experiments. We provide further discussion and experiments on online scheduling for dynamic workloads, as well as on the trade-offs between cost-efficiency and request latency, in Appendix M.
@@ 6 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgment
This work is supported by the HKUST startup grant R9895 from CSE; RGC-ECS project 26218024; RGC-NSFC project CRS_HKUST601/24. This work was supported as part of the Swiss AI Initiative by a grant from the Swiss National Supercomputing Centre (CSCS) on Alps.
@@ Appendix E Extend to Multiple LLM serving
s.t. $\forall m:\begin{cases}\sum_{c\in\mathcal{C}_{m}}x_{c,w,m}=1,\,\forall\,w\in W_{m},\\[4.0pt] \sum_{w\in W_{m}}\frac{x_{c,w,m}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\cdot f_{w,m}}}{y_{c,m}\cdot h_{c,w,m}}\leq T,\,\forall\,c\in\mathcal{C}_{m},\\[4.0pt] x_{c,w,m}\leq y_{c,m},\,\forall\,c\in\mathcal{C}_{m},\,\forall\,w\in W_{m},\end{cases}$ (9)
@@ Appendix I Workload Type Ratios for Each Trace
Table 6: Performance of different configurations.
Diff Configs Real Estimated
H100 (2,4) 0.56 req/s 0.60 req/s
H100 (4,2) 0.44 req/s 0.47 req/s
H100 (4,2) (cross machine) 0.42 req/s 0.44 req/s
L40 (2,4) 0.42 req/s 0.46 req/s
L40 (4,2) 0.21 req/s 0.22 req/s
L40 (4,2) (cross machine) 0.18 req/s 0.19 req/s
H100+A100 (4,2) (cross machine) 0.48 req/s 0.52 req/s
Appendix L One-Time Profiling
We employ a one-time profiling strategy that captures the following components. This approach is referred to the profiling method used in Vidur (Agrawal et al., 2024a): (1) Inference prefilling latency: We profile the latency for a single transformer layer across varying TP degrees, different workload types, and various GPU types; (2) inference decoding latency: We profile the decoding latency for a single transformer layer under similar variations in TP degrees, workload types, and GPU types; (3) pipeline communication latency: We measure the communication latency between different GPUs across various workload types. Using these measurements, the per-request latency for any configuration is estimated by combining the TP costs (both communication and computation) of all layers—which may be served by different GPUs and at varying TP degrees—with the PP communication cost. Note that our heuristics, as discussed in Section 4.3 and Appendix D, largely reduce the profiling space, e.g., TP is employed only intra-machine. When estimating throughput, the prefill and decoding phases are treated separately: (1) The prefill phase is compute-bound, and its batched processing capacity is determined by the sum of the individual latencies; (2) the decoding phase is memory-bound, with its batched processing capability defined by a single latency value. This distinction has been validated in several studies (Zhong et al., 2024; Patel et al., 2024).
Table 6 demonstrates examples of our cost estimation under a long-input, short-output workload (i.e., workload 1 in Figure 4). In Table 6, the notation (2,4) indicates that the TP degree is 2 and the PP degree is 4. The estimation errors range from 4% to 7%. Although the estimations are not perfectly accurate, they are sufficiently reliable for selecting the optimal configurations.
Appendix M Discussion
Online replanning. Online scheduling for dynamic workloads is orthogonal to the approach presented in this work. However, accommodating dynamic workloads could be achieved through the implementation of a replanning mechanism analogous to the one proposed in DistServe (Zhong et al., 2024). Concretely, the system could (1) monitor the real-time composition of incoming workloads, (2) track GPU resource availability within the cloud environment, and (3) upon detecting a significant shift in workload distribution, (e.g., an increase in the proportion of certain workload types) the scheduling algorithm could be executed again, incorporating recent historical data to produce an updated serving plan.
Table 7: Performance changes in workload and GPU drop.
Workload Change GPU Drop
Before 26.89 req/s 26.89 req/s
After 23.70 req/s (13$\downarrow$) 20.80 req/s (29$\downarrow$)
Replanning 29.61 req/s (25$\uparrow$) 22.85 req/s (10$\uparrow$)
As shown in Table 7, we test the workload surge in short output requests in AzureTrace dataset (i.e., trace 2) with a price budget of \$30 per hour. Before surge, the optimal GPU composition is {20%, 65%, 15%} for datacenter, workstation, and consumer GPUs, achieving 26.89 req/s. After workload change, the throughput degrades to 23.7 req/s. In this case, replanning (shifting allocation to {63%, 23%, 14%}) boosts throughput to 29.61 req/s. We also test the case when GPU drop happens (4 H100s down), the throughput falls from 26.89 to 20.80 req/s. In this case, replanning raises throughput to 22.85 req/s.
Trade-offs between prioritizing cost-efficiency and request latency. Prioritizing cost-efficiency typically involves using fewer resources (i.e., lower budgets), which can lead to slightly higher response latencies. In contrast, prioritizing latency often requires utilizing more resources (i.e., incurring higher costs). We acknowledge that optimizing for cost efficiency may result in a slight increase in latency. However, inference tasks typically do not require extremely low latency; meeting a predefined latency threshold is usually sufficient. In resource-limited scenarios, where systems are naturally under-provisioned, emphasizing throughput can also indirectly improve latency by reducing queuing delays. Our experimental results in Figure 6 demonstrate that our method achieved the lowest P99 latency among all compared baselines.